# Towards Efficient Transferable Preemptive Adversarial Defense

### v3

@@ III-C2 Preliminary - Backward Propagation
where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size. The final output of Equation (8) is adversarial example $x_{t}^{a}$. Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
@@
III-C3 Combining Forward and Backward Propagation Learning
We define forward propagation learning as a direct enhancement by moving away from the decision boundary in accordance with the learning rate ($x_{0}\rightarrow x_{1}$ in Figure 2). Therefore, it first reduces $\ell$ via
@@ III-C3 Combining Forward and Backward Propagation Learning
which is the same as for Equation (9). Therefore, the above derivations can be summarized as
@@
where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation (12) indicates that both forward and backward propagation make the image more robust but with different learning directions for CE loss $\ell$.
@@
III-C4 Averaging $N$ Fast Gradient Sign Method (FGSM) Examples
@@
Forward and backward propagation both require a backbone optimization algorithm, i.e., an adversarial attack algorithm. For example, The Bi-Level optimization [10] used PGD [17] for backward propagation. However, PGD has to run many sequential iterations to compute superior adversarial perturbations, which greatly increases the time cost. With the Fast Preemption strategy, FGSM [1] is used to speed up the computation of adversarial perturbations.
@@
Since FGSM has difficulty obtaining optimal perturbations, with the Fast Preemption strategy, it is run $N$ times with random initialization to produce $N$ FGSM examples, The perturbations are then averaged to obtain the final output. We demonstrate that averaging the perturbations for PGD examples negligibly affects performance (see Section VII-D), whereas with the Fast Preemption strategy, FGSM optimization is effectively improved while efficiency is preserved.
@@
Furthermore, averaging perturbations in the $N$ FGSM examples is faster than processing a single example with iterative PGD. To generate a single RE, iterative PGD has to run each iteration sequentially from a single input, whereas $N$ FGSM examples can be produced by repeating the single input $N$ times and processing the results in parallel in a single iteration. The improvement attained by averaging the perturbations in $N$ FGSM examples is demonstrated in Table VIII.
@@
III-C5 Algorithm
The complete Fast Preemption algorithm is presented in Algorithm 1. With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel.
IV Adaptive Reversion Attack:
Preemptive Reversion
IV-A Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Figure 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Figure 3. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
The complete algorithm is presented in Algorithm 2.
@@ IV-A Algorithm
IV-B Attack Scenarios
  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (due to infinite selections for users). The major challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our proposed Fast Preemption defense, but it is feasible only in the white-box scenario. Without full knowledge of the defense, Fast Preemption is reliable.
IV-C Evaluation Protocol
If a reversion attack succeeds, the improved clean accuracy should drop. However, image distortion can create a similar effect. Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack. To address this issue, we proposed a more comprehensive protocol that can distinguish successful reversion from distortion. First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy. In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded. Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
TABLE I: 1,000 original CIFAR-10 [15] images show similar accuracy with that of full test set against AutoAttack (Standard) [18] with training-time defense [7] @ $L_{\infty}\ \epsilon=8/255$.
     Number of Test Samples      Clean (%)      Robust (%)
     10,000 (full)      93.27      71.07
     1,000 (ours)      93.5      71.4
     100      95      68
V Experiments
V-A Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [15] (size of 32$\times$32) and ImageNet [16] (cropped to 224$\times$224 to match the requirements of the classifier and backbone model). CIFAR-10: We conducted ablation studies using 1,000 samples and then tested our defense and compared the results with benchmarks using another 1,000 samples (resembling the validation and test datasets). ResNeXt29-32x4d [37] and PreActResNet-18 [5] were selected as the classifier and backbone model, respectively. Because Fast Preemption does not require model training, 1,000 samples provided a satisfactory diversity and distribution for testing, as demonstrated by the results listed in Table I. ImageNet: We evaluated Fast Preemption and benchmarks using the development dataset from the NIPS 2017 competition, including 1,000 carefully selected ImageNet samples. DeiT Base [38] and ConvNeXt-L+ConvStem [14] were selected as the classifier and backbone model, respectively. The model architectures are listed in Table II.
TABLE II: Model architectures.
@@ V-A Datasets, Classifier, and Backbone Models
V-B Benchmark Adversarial Defenses
Training-time defenses: We compared Fast Preemption with PreActResNet-18 [5], XCiT-L12 [6], and RaWideResNet-70-16 [7] on the CIFAR-10 dataset. We further compared it with Swin-L [8] on the ImageNet dataset. RaWideResNet-70-16 [7] and Swin-L [8] are the SOTA training-time defenses for each dataset, respectively. As listed in Table II, the model architectures of the SOTA training-time defenses differ from the backbone models, so our experiments are a reasonable way to evaluate transferability.
Test-time defenses: We used DiffPure [9] as the benchmark test-time defense. The inputs (regardless of original or malicious) were purified by DiffPure and then fed into RaWideResNet-70-16[7] or Swin-L [8] in accordance with the dataset used.
Preemptive defenses. We compared Fast Preemption with two benchmark preemptive defenses: the SOTA preemptive defense, Bi-Level [10], and the latest preemptive defense, $A^{5}$ [11]. We used the best defense recipe, “RC (robustifier and classifier),” given by Frosio and Kautz [11].
V-C Attacks
Adversarial attacks. We evaluated defenses against two adversarial attacks: PGD [17] and AutoAttack [18]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [39] and Square Attack [40]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [41]).
Reversion attacks. We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [9]).
V-D Evaluation metrics
We used both clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective. Here, accuracy refers to the ratio of predictions that match their ground-truth labels. Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs. Additionally, to capture the image quality, the structural similarity index measure (SSIM) [42] was computed with respect to the original images. A larger SSIM indicates a more subtle change.
V-E Implementation
VI Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Table III.
TABLE III: Experiment Settings.
@@ VI Performance
3
Ours Same with our defense
@@
3
Bi-Level $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
$\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
3
@@
3
Ours $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$
$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
@@
TABLE IV: End-to-end performance against AutoAttack [18] and comparison with benchmark adversarial defenses.
CIFAR-10 Training-Time RaWideResNet-70-16 [7] 0 93.3 72.8 69.3 73.4
8
Test-Time DiffPure [9] 0.3 91.4 75.8 83.7 76.2
8
Preemptive $A^{5}$ [11] 0.01 93.2 72.4 68.1 73.0
8
Bi-Level [10] 28 92.0 80.1 78.0 80.2
8
Fast Preemption (ours) 0.04 95.3 85.6 84.5 85.7
ImageNet Training-Time Swin-L [8] 0 93.0 67.0 59.5 67.5
8
Test-Time DiffPure [9] 1.2 88.8 73.7 79.4 73.8
8
Preemptive Bi-Level [10] 43 93.6 71.5 62.8 72.3
8
Fast Preemption (ours) 1.3 95.1 77.1 67.4 78
8
Fast Preemption (ours) + DiffPure [9] 2.5 93.3 82.9 87.6 83.4
TABLE V: Preemptive defenses resist reversion attacks.
9
93.5 Bi-Level Bi-Level Bi-Level 91.9 90.5 90.8 93.1 90.7
9
Bi-Level PreActResNet-18 Bi-Level 91.2 90.6 90.7 93.4 91.0
9
Bi-Level PreActResNet-18 Ours 91.3 90.5 90.5 93.4 90.6
9
ResNeXt29-32x4d PreActResNet-18 Ours 95.5 94.4 94.8 94.0 94.6
VI-A Robustness against Adversarial Attacks
We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [18] and victim models [7, 8]. Table IV compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [18]. Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive. For example, in comparison with the least time costly method RaWideResNet-70-16 [7] with the CIFAR-10 dataset, Fast Preemption had 2% and 12.8% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation. Compared with the SOTA preemptive defense, Bi-Level [10], which had the second best accuracy, Fast Preemption had 3.3% and 5.5% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster. Note that preemptive defenses can work together with training-time and test-time defenses. For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last row of Table IV.
Figure 4 illustrates the results of qualitative analysis. A visual comparison of the original image and RE (average SSIM 0.962) revealed that the perturbations generated by Fast Preemption are unnoticeable. Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones. Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [45]:
$\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
$\delta_{Gray}=$ $\ 0.299\cdot\delta_{RGB}.R+$
$\ 0.587\cdot\delta_{RGB}.G+0.114\cdot\delta_{RGB}.B,$
VI-B Resistance against Reversion Attacks
We conducted reversion attacks to evaluate whether preemptive defenses are reversible. Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense. Preemptive Reversion was evaluated in both white- and black-box scenarios (see Section IV-B). The results are summarized in Table V. The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense. The other reversion attacks failed because they further reduced clean accuracy (due to distortion). Nevertheless, Preemptive Reversion is only feasible in the white-box setting. Without full knowledge of the defense, Fast Preemption is irreversible. Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails. However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
Moreover, the results in the fourth row of Table V, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive. The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion. However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion. Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
VII Ablation Studies
We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset. The attacks were PGD [17] and AutoAttack (Standard) [18] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. As illustrated in Figure 6, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [10] in terms of clean and robust accuracies and time cost. These studies and results were aimed at identifying the optimal strategies.
VII-A Classifier
In all proactive defenses, correct labeling guarantees that optimization moves in the correct direction. As demonstrated in Table VI, a classifier with higher accuracy leads to better performance. As this classifier is not involved in the Fore-Back learning, the time cost of labeling inputs can be ignored compared with that of Fore-Back learning. Note that the use of ground truths is ideal, but this requires manual labeling, which is impractical for efficient user interaction. Therefore, we used the best pretrained classifier in the Fast Preemption framework, one that differs from the backbone model.
TABLE VI: Performance with different classifiers.
@@ VII-A Classifier
  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [10]. The transferable defenses were all evaluated using RaWideResNet-70-16 [7] as the victim model.
VII-B Backbone Model
The backbone model determines the defense time cost and performance. As illustrated in Figure 7, different backbone models learn different features. To evaluate the effect of the backbone model, we also evaluated white-box defenses, for which the backbone model is the same as the victim model, i.e., paired modules, as stated in Section I, although we believe that this setting is not secure due to potential privacy leakage due to the existence of a backdoor [13].

### v4

@@ 3.3.2 Preliminary - Backward Propagation
where $\epsilon$ is the setting of the maximum distance and $\alpha$ is the step size. The final output of Eq. 8 is AE $x_{t}^{a}$. Next, backward propagation subtracts perturbations $x_{t}^{a}-x_{t}$ from $x_{t}$ in accordance with learning rate $\eta$, thereby computing the next robust example, $x_{t+1}$:
@@
3.3.3 Combining Forward and Backward Propagation
We define forward propagation learning as a direct enhancement by moving away from the decision boundary in accordance with the learning rate ($x_{0}\rightarrow x_{1}$ in Fig. 2). Therefore, it first reduces $\ell$ via
@@ 3.3.3 Combining Forward and Backward Propagation
which is the same as for Eq. 9. Therefore, the above derivations can be summarized as
@@
where $x_{*}$ denotes $x_{t,i+1}$ or $x_{t+1}$. Equation 12 indicates that both forward and backward propagation make inputs more robust but with different learning directions for loss $\ell$.
@@
3.3.4 Averaging $N$ Examples
@@
Forward and backward propagation both require a backbone optimization algorithm, i.e., an adversarial attack algorithm. For example, The Bi-Level optimization [29] used PGD [28] for backward propagation. However, PGD has to run many sequential iterations to compute superior adversarial perturbations, which greatly increases the time cost. With the Fast Preemption strategy, Fast Gradient Sign Method (FGSM) [16] is used to speed up the computation of adversarial perturbations.
@@
Since FGSM has difficulty obtaining optimal perturbations, with the Fast Preemption strategy, it is run $N$ times with random initialization to produce $N$ FGSM examples, The perturbations are then averaged to obtain the final output. We demonstrate that averaging the perturbations for PGD examples negligibly affects performance, whereas with the Fast Preemption strategy, FGSM optimization is effectively improved while efficiency is preserved.
@@
Furthermore, averaging perturbations in the $N$ FGSM examples is faster than processing a single example with iterative PGD. To generate a single RE, iterative PGD has to run each iteration sequentially from a single input, whereas $N$ FGSM examples can be produced by repeating the single input $N$ times and processing the results in parallel.
4 Adaptive Attack: Preemptive Reversion
4.1 Algorithm
When running preemptive defenses sequentially for the same input and settings, the perturbations are mostly the same for two rounds (see Fig. 5). On the basis of this finding, we devised the first effective white-box adaptive reversion attack, i.e., Preemptive Reversion, as illustrated in Fig. 3. The complete algorithm is presented in Algorithm 2 of Appendix B. Using RE $x^{r}$ generated by a preemptive defense method, the Preemptive Reversion algorithm first runs a secondary preemptive defense upon $x^{r}$ to obtain $(x^{r})^{r}$. Then, the Preemptive Reversion algorithm reverses the secondary perturbations from $x^{r}$ to obtain $\tilde{x}$, which is the prediction of the original input $x$:
$\tilde{x}=x^{r}-((x^{r})^{r}-x^{r}).$ (13)
4.2 Attack Scenarios
  • Black-Box: The attacker knows the defense algorithm (typically public) but not the backbone model (infinite selections for users). The challenge for the attacker in this scenario is that the crucial features learned in the secondary defense may differ from those learned by the user.
Our proposed Preemptive Reversion attack is the first effective reversion attack for evaluating the reversibility of preemptive defenses, including our Fast Preemption defense, but it is feasible only in the white-box scenario. Without full knowledge of the defense, Fast Preemption is reliable.
4.3 Evaluation Protocol
If a reversion attack succeeds, the improved clean accuracy should drop. However, image distortion can create a similar effect. Consequently, a drop in clean accuracy does not necessarily indicate the success of a reversion attack. To address this issue, we proposed a comprehensive protocol that can distinguish successful reversion from distortion. First, we purposely give incorrect labels to partial inputs to guarantee that the clean accuracy after being defended is lower than the original accuracy. In this situation, if the clean accuracy approaches close to the original accuracy, the reversion attack succeeded. Otherwise, the reversion attack failed, and the deterioration was caused by distortion.
5 Experiments
5.1 Datasets, Classifier, and Backbone Models
We conducted experiments on two datasets: CIFAR-10 [24] and ImageNet [1]. For CIFAR-10, ResNeXt29-32x4d [19] and PreActResNet-18 [17] were selected as the classifier and backbone model, respectively. For ImageNet, DeiT Base [41] and ConvNeXt-L+ConvStem [40] were selected as the classifier and backbone model, respectively. The model architectures are listed in Tab. 3 of Appendix C.
5.2 Attacks and Defenses
Adversarial attacks. We evaluated defenses against two adversarial attacks: PGD [28] and AutoAttack [9]. For AutoAttack, we used both its “Standard” version (combining APGD-CE, APGD-DLR, and FAB [8] and Square Attack [3]) and its “Random” version (combining APGD-CE and APGD-DLR, with expectation over time (EOT) equal to 20 [4]). Reversion attacks. We ran reversion attacks to evaluate: (i) whether the protected samples induced by preemptive defenses cannot be easily and precisely reversed back to their original unprotected status, which we define as “irreversible” in this paper; (ii) whether the proposed adaptive reversion attack, Preemptive Reversion, is effective. Therefore, apart from the specific Preemptive Reversion, we also ran reversion attacks using blind distortion (JPEG compression) and adversarial purification (DiffPure [31]).
Training-time defenses: We compared Fast Preemption with WideResNet-94-16 [5] and Swin-L [26] on the CIFAR-10 and ImageNet datasets, respectively. The model architectures of these defenses differ from the backbone models. Test-time defenses: We used DiffPure [31] as the benchmark test-time defense. The inputs were purified by DiffPure and then fed into WideResNet-94-16 [5] or Swin-L [26] in accordance with the dataset used. Preemptive defenses. We compared Fast Preemption with the SOTA preemptive defense, Bi-Level [29], and the latest preemptive defense, $A^{5}$ [15].
5.3 Evaluation metrics
We used clean accuracy and robust accuracy for performance evaluation, where higher accuracy is better from the defense perspective. Here, accuracy refers to the ratio of predictions that match their ground-truth labels. Clean accuracy was computed for benign images (original and REs) while the robust accuracy was computed for AEs.
Table 1: End-to-end performance against AutoAttack [9] and comparison with benchmark adversarial defenses.
CIFAR-10 Training-Time WideResNet-94-16 [5] 0 93.4 74.6 69.6 74.8
8
Test-Time DiffPure [31] 0.3 91.4 78.1 84.9 78.4
8
Preemptive $A^{5}$ [15] 0.01 93.5 73.8 69.0 74.3
Bi-Level [29] 28 92.5 81.4 78.6 81.6
Fast Preemption (ours) 0.04 95.0 86.5 84.3 86.7
8
Fast Preemption (ours) + DiffPure [31] 0.34 94.2 89.1 92.8 89.7
ImageNet Training-Time Swin-L [26] 0 93.0 67.0 59.5 67.5
8
Test-Time DiffPure [31] 1.2 88.8 73.7 79.4 73.8
8
Preemptive Bi-Level [29] 43 93.6 71.5 62.8 72.3
Fast Preemption (ours) 1.3 95.1 77.1 67.4 78
8
Fast Preemption (ours) + DiffPure [31] 2.5 93.3 82.9 87.6 83.4
  • Bold indicates the best performance.
6 Performance
In this section, we present the end-to-end evaluation results and comparisons with benchmarks. The results were obtained in the settings listed in Tab. 4 of Appendix D. The time cost was recorded with an NVIDIA A100 40GB GPU, as described in Appendix E.
Table 2: Preemptive defenses resist reversion attacks.
93.5 Bi-Level Bi-Level Bi-Level 91.9 90.5 90.8 93.1 90.7
Bi-Level PreActResNet-18 Bi-Level 91.2 90.6 90.7 93.4 91.0
Bi-Level PreActResNet-18 Ours 91.3 90.5 90.5 93.4 90.6
ResNeXt29-32x4d PreActResNet-18 Ours 95.5 94.4 94.8 94.0 94.6
6.1 Robustness against Adversarial Attacks
We evaluated transferable defense performance for the situation in which the defense has no knowledge of the AutoAttack [9] and victim models [5, 26]. Table 1 compares the performance of Fast Preemption with that of the benchmark defenses against AutoAttack [9]. Fast Preemption performance was better than that of the benchmarks in terms of clean and robust accuracies, whereas its time cost was competitive. For example, in comparison with the least time costly method WideResNet-94-16 [5] with the CIFAR-10 dataset, Fast Preemption had 1.6% and 11.9% better clean and robust accuracies (Standard $L_{\infty}$), respectively, with 0.04 more seconds per generation. Compared with the SOTA preemptive defense, Bi-Level [29], which had the second best accuracy, Fast Preemption had 2.5% and 5.1% better clean and robust accuracies (Standard $L_{\infty}$), respectively and was 700 times faster. Note that preemptive defenses can work together with training-time and test-time defenses. For example, when combined with a test-time defense, the performance of Fast Preemption was remarkably improved, as reflected in the last rows of CIFAR-10 and ImageNet in Tab. 1, respectively.
Figure 4 illustrates the results of qualitative analysis. A visual comparison of the original image and RE revealed that the perturbations generated by Fast Preemption are unnoticeable. Moreover, the “Success” and “Failure” cases indicate that preemptive defense is more likely to succeed when the perturbations are similar to the attack ones. Note that the perturbation illustration $\delta_{Gray}$ was drawn by first normalizing the true perturbations $\delta$ from $[-\epsilon,+\epsilon]$ to $[0,1]$ to produce $\delta_{RGB}$ and then transforming $\delta_{RGB}$ into grayscale using the National Television Standards Committee (NTSC) formula [27]:
$\delta_{RGB}=$ $\ \frac{\delta+\epsilon}{2\cdot\epsilon},$ (14)
$\delta_{Gray}=$ $\ 0.299\cdot\delta_{RGB}.R+$
$\ 0.587\cdot\delta_{RGB}.G+0.114\cdot\delta_{RGB}.B,$
6.2 Resistance against Reversion Attacks
We conducted reversion attacks to evaluate whether preemptive defenses are reversible. Apart from the proposed Preemptive Reversion attack, we used JPEG compression and adversarial purification because they require no knowledge of the defense. Preemptive Reversion was evaluated in both white- and black-box scenarios (see Sec. 4.2). The results are summarized in Tab. 2. The results shown in red approach the original accuracy, demonstrating that only our reversion attack was able to reverse the preemptive defense. The other reversion attacks failed because they further reduced clean accuracy (due to distortion). Nevertheless, Preemptive Reversion is only feasible in the white-box setting. Without full knowledge of the defense, Fast Preemption is irreversible. Figure 5 further supports our claim that the white-box reversion successfully erases most protective perturbations while the black-box reversion fails. However, the white-box reversion cannot fully remove the protective perturbations due to the randomization in Fast Preemption (i.e., randomly initialized FGSM).
Moreover, the results in the fourth row of Tab. 2, which were obtained without using our evaluation protocol, demonstrate that our protocol for evaluating reversion attacks is more comprehensive. The results shown in the other rows demonstrate that our Preemptive Reversion in the white-box setting is the only feasible reversion attack as it is the only one for which clean accuracy was higher after reversion. However, the lower numbers in the fourth row indicate that the decrease in accuracy with Preemptive Reversion in the white-box setting was due to successful reversion, whereas the other decreases were due to distortion. Therefore, if only the results in the fourth row are considered, a decrease in accuracy does not necessarily indicate the effectiveness of a reversion attack.
7 Ablation Studies
We conducted ablation studies on the Fast Preemption framework and its learning algorithm using the CIFAR-10 dataset. We used AutoAttack (Standard) [9] against a black-box target model RaWideResNet-70-16 [32], employing $L_{\infty}$-norm perturbations with $\epsilon=8/255$. As demonstrated in Fig. 6, Fast Preemption significantly outperforms the SOTA Bi-Level method [29], increasing clean accuracy from 92% to 95.3%, improving robust accuracy from 80.1% to 85.6%, and reducing computation time from 28 seconds to just 0.04 seconds. These studies were conducted to identify the optimal configurations, with more extensive ablation results presented in Appendix F.
8 Conclusion
We have presented an efficient transferable preemptive adversarial defense that we call Fast Preemption. We demonstrated that with its framework and Fore-Back cascade learning algorithm, Fast Preemption offers transferable protection against adversarial attacks among unknown victim models without requiring ground truths. In terms of the time cost, clean accuracy, and robust accuracy, Fast Preemption outperformed benchmark training-time, test-time, and preemptive defenses. We also demonstrated that the protection from Fast Preemption is irreversible unless the backbone model, algorithm, and settings of the defense are all compromised. This paper lays the theoretical foundation for a preemptive defense framework that can shape future research on each component’s optimization instead of end-to-end application in the relevant field.
Acknowledgments
This work was supported by JSPS KAKENHI Grants JP21H04907 and JP24H00732, by JST CREST Grants JPMJCR18A6 and JPMJCR20D3 including AIP challenge program, by JST AIP Acceleration Grant JPMJCR24U3, by JST K Program Grant JPMJKP24C2 Japan, and by the project for the development and demonstration of countermeasures against disinformation and misinformation on the Internet with the Ministry of Internal Affairs and Communications of Japan.
Supplementary Material
This supplementary material provides an in-depth discussion of individual related works (Appendix A), the algorithms underpinning the proposed adversarial defense and adaptive reversion attack (Appendix B), model architectures (Appendix C), and the experimental configurations for both attacks and defenses (Appendix D). Additionally, it details the implementation specifics (Appendix E) and presents comprehensive ablation study results Appendix F, covering aspects such as the classifier, backbone models, the Forward-Backward Cascade Learning algorithm, and the strategy of averaging multiple Fast Gradient Sign Method (FGSM) examples.
Appendix A Related Work
Training-time defenses produce robust deep learning models that may correctly classify adversarial examples (AEs) as their ground truth labels. The earliest mechanisms, e.g., adversarial training [16], incorporated various types of AEs along with their ground-truth labels in the training dataset. Zhang et al. [46] proposed TRADES, which introduces a theoretically grounded approach to balance robustness and accuracy in adversarial training by incorporating a custom loss function that controls the trade-off between clean accuracy and adversarial robustness. To improve the efficiency of training on large-scale datasets, Shafahi et al. [38] devised accelerated adversarial training in which backward-pass computations are reused. This entails curating a large number of specific real-world samples; Sehwag et al. [37] circumvented this problem by using additional data from proxy distributions learned by advanced generative models for adversarial training. Gowal et al. [17] demonstrated that generative models trained solely on the original dataset can be leveraged to artificially increase the training dataset’s size and improve adversarial robustness against $L_{p}$-norm perturbations. Debenedetti et al. [11] proposed using a custom recipe of vision transformers for adversarial training in which weaker data augmentation is used with warm-up in the augmentation intensity. Inspired by denoising diffusion models [21], Wang et al. [45] conducted adversarial training using data generated by an elucidating diffusion model [22]. Dong et al. [12] introduced robust examples (REs) generated by a preemptive defense into the adversarial training and demonstrated alleviation of clean accuracy deterioration. Peng et al. [32] devised a robust model which follows a suite of generalizable robust architectural design principles: (i) an optimal range for depth and width configurations; (ii) a preference for a convolutional rather than patchify stem stage, and (iii) a robust residual block design featuring squeeze and excitation blocks and nonparametric, smooth activation functions. Bartoldson et al. [5] developed the current state-of-the-art (SOTA) robust model by scaling laws for adversarial training to assess robustness in image classifiers, revealing that even with significant computational resources, robustness asymptotes below human-level performance due to flaws in attack formulations generating invalid data, suggesting future efforts should address both computational efficiency and image validity.
Training-time defenses have two shortcomings. First, they require vast amounts of data for training, and the data quality and diversity are crucial to robustness. Second, training-time defenses typically reduce clean accuracy in favor of robust accuracy, even the one proposed by Dong et al. [12], which was designed to circumvent this problem.
Test-time defenses are divided to detection and purification. Detection models are typically trained on AEs to enable the model to distinguish them from benign inputs. Among recent studies, Cohen et al. [7] devised a k-nearest neighbor model for ranking the most supportive training samples for any given validation example in training a detector. Zuo and Zeng [47] proposed using a binary classifier, erase-and-restore, that uses the intriguing sensitivity of $L_{2}$-norm attacks by randomly erasing certain pixels in an $L_{2}$-norm AE and then restoring them with an inpainting technique. Liang et al. [25] detected AEs by comparing the classification results of a given sample and a denoised version that was generated through adaptive noise reduction combining scalar quantization and a smoothing spatial filter. Most recently, Hickling et al. [20] devised a detector using explainable deep reinforcement learning for uncrewed aerial vehicle agents. However, detection of unseen attacks is challenging for all existing detectors because they are trained on AEs from known attacks [6]. Additionally, detection always introduces errors due to false classification of benign images as AEs, thus reducing clean accuracy.
In contrast, adversarial purification aims to purify AEs by eliminating or invalidating perturbations. Earlier studies used image distortion for purification. For instance, Wang et al. [44] distorted images by randomly nullifying pixels. Guo et al. [18] preprocessed inputs with a nondifferentiable transformation. Raff et al. [33] combined a large number of individually weak defenses into a single randomized transformation to build a strong defense. Lately, more practical approaches generate benign replacements for inputs regardless of whether they are malicious. Ren et al. [34] trained a generative model using autoencoder technology [23]. They claimed the existence of an immune space, in which perturbations have fewer adverse effects, and they inactivated adversarial perturbations by restricting them to that subspace. Nie et al. [31] devised the DiffPure benchmark, which uses denoising diffusion models [21]. DiffPure replaces perturbations with Gaussian noise and then restores the clean image by reversing the diffusion process. The major limitation of purification is that the difference between the original and generated examples degrades clean accuracy, and the possibility of being counterattacked [42].
@@ Appendix A Related Work
Appendix B Algorithms of Fast Preemption and
Preemptive Reversion
The complete defense algorithm of Fast Preemption is presented in Algorithm 1. With PyTorch implementation, Steps 4 to 8 and Steps 15 to 19 can be respectively processed in parallel. The complete adaptive attack algorithm of Preemptive Reversion is presented in Algorithm 2.
Appendix C Model Architectures
We evaluated attacks and defenses using models with various architectures, as listed in Tab. 3.
Table 3: Model architectures.
@@ Appendix C Model Architectures
Table 4: Experiment Settings.
@@
        Fast Preemption (ours)         Same with our defense
@@
        Bi-Level         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$, $\eta^{3}=0.1$, $T_{\leftarrow}=100$, $N=1$
@@
        Fast Preemption (ours)         $L_{\infty}\ \epsilon=\frac{8}{255}^{1}/\frac{4}{255}^{2}$,$\eta^{3}=0.7$, $T_{\rightarrow}=1$, $T_{\leftarrow}=2$, $N=20$
@@
Table 5: Performance with different classifiers.
@@
  • To evaluate the performance of diverse classifiers in our framework, we used the same backbone model and 100 steps backward propagation as Bi-Level [29]. The transferable defenses were all evaluated using RaWideResNet-70-16 [32] as the victim model. Bold indicates the best performance.
Appendix D Experiment Settings
The full experimental settings used for evaluations are listed in Tab. 4.
Appendix E Implementation
Appendix F Ablation Studies
We conducted ablation studies on the Fast Preemption framework and learning algorithm using the CIFAR-10 dataset. The attacks were PGD [28] and AutoAttack (Standard) [9] with $L_{\infty}$-norm perturbations and $\epsilon=8/255$. As illustrated in Fig. 6 of Sec. 7, Fast Preemption with the proposed framework and learning algorithm outperformed the SOTA benchmark Bi-Level [29] in terms of clean and robust accuracies and time cost. These studies and results were aimed at identifying the optimal strategies.